OpenStudy (alayrhi):

What is 5/6 + (-1/2)

OpenStudy (steve816):

First of all, a + and a - will make it negative, so you are doing \[\Large \frac{ 5 }{ 6 }-\frac{ 1 }{ 2 }\]

OpenStudy (steve816):

Next, in order to subtract the fractions, you need a common denominator ya?

OpenStudy (alayrhi):

1/2 = 3/6

OpenStudy (steve816):

That's right! Now can you do the subtraction?

OpenStudy (alayrhi):

2/6

OpenStudy (steve816):

That is right! And you can reduce it even further to 1/3

OpenStudy (alayrhi):

so the answer is...1/3!?

OpenStudy (steve816):

Yup!

OpenStudy (alayrhi):

ahh okay thanks ^_^

OpenStudy (steve816):

You're welcome :)